# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

### v1

@@ 2 Incremental Gradient Descent and Shuffled SGD
We begin the analysis by deriving a bound on the gap with respect to an arbitrary but fixed reference point ${\bm{z}},$ as summarized in the following lemma. This stands in contrast to arguments deriving bounds on the average iterate, which take ${\bm{z}}={\bm{x}}_{*}.$ While this may seem like a minor difference, it affects the analysis non-trivially: a direct extension of prior arguments would require replacing 4 – which imposes a bound on $\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{*})\|^{2}$ – with a bound on $\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})\|^{2}$ for an arbitrary ${\bm{z}}$, which would be a much stronger requirement.

### v2

@@ 2 Last Iterate Convergence of Incremental Gradient Descent
We begin the analysis by deriving a bound on the gap with respect to an arbitrary but fixed reference point ${\bm{z}},$ as summarized in the following lemma with its proof in Appendix A. This stands in contrast to arguments deriving bounds on the average iterate, which take ${\bm{z}}={\bm{x}}_{*}.$ While this may seem like a minor difference, it affects the analysis non-trivially: a direct extension of prior arguments would require replacing Assumption 4 – which imposes a bound on $\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{*})\|^{2}$ – with a bound on $\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})\|^{2}$ for an arbitrary ${\bm{z}}$, which would be a much stronger requirement.